Suppose that April has access to an investment that will pay 10% interest compounded continuously. Which is better : to be given $1000 now so that she can take advantage of this investment opportunity or to be given $1325 after 3 years?
Answer by Boreal(15235)

P=Poe^rt
=1000*e^.10*3=10100*e^0.30=$1349.86
This will allow both taking advantage of the investment or, if nothing is done, will be more than it will be to receive $1325 in 3 years.